### v2

@@ 3 Discussion
We hope that by making our data and code publicly available, we will help the field develop a data-driven, impartial understanding of the current opportunities and challenges presented by digital twins.
@@
A Mega-study of Digital Twins Reveals Strengths, Weaknesses and Opportunities for Further Improvement
@@
appendix.a.Aappendix.a.Bsubsection.a.B.1subsection.a.B.2subsection.a.B.3appendix.a.Csubsection.a.C.1subsection.a.C.2subsection.a.C.3appendix.a.Dappendix.a.Eappendix.a.Fappendix.a.Gsubsection.a.G.1subsection.a.G.2appendix.a.Happendix.a.Iappendix.a.Jsubsection.a.J.1subsubsection.a.J.1.1subsubsection.a.J.1.2subsubsection.a.J.1.3subsubsection.a.J.1.4subsubsection.a.J.1.5subsection.a.J.2subsubsection.a.J.2.1subsubsection.a.J.2.2subsubsection.a.J.2.3subsubsection.a.J.2.4subsubsection.a.J.2.5subsection.a.J.3subsubsection.a.J.3.1subsubsection.a.J.3.2subsubsection.a.J.3.3subsubsection.a.J.3.4subsubsection.a.J.3.5subsection.a.J.4subsubsection.a.J.4.1subsubsection.a.J.4.2subsubsection.a.J.4.3subsubsection.a.J.4.4subsubsection.a.J.4.5subsection.a.J.5subsubsection.a.J.5.1subsubsection.a.J.5.2subsubsection.a.J.5.3subsubsection.a.J.5.4subsubsection.a.J.5.5subsection.a.J.6subsubsection.a.J.6.1subsubsection.a.J.6.2subsubsection.a.J.6.3subsubsection.a.J.6.4subsubsection.a.J.6.5subsection.a.J.7subsubsection.a.J.7.1subsubsection.a.J.7.2subsubsection.a.J.7.3subsubsection.a.J.7.4subsection.a.J.8subsubsection.a.J.8.1subsubsection.a.J.8.2subsubsection.a.J.8.3subsubsection.a.J.8.4subsection.a.J.9subsubsection.a.J.9.1subsubsection.a.J.9.2subsubsection.a.J.9.3subsubsection.a.J.9.4subsection.a.J.10subsubsection.a.J.10.1subsubsection.a.J.10.2subsubsection.a.J.10.3subsubsection.a.J.10.4subsection.a.J.11subsubsection.a.J.11.1subsubsection.a.J.11.2subsubsection.a.J.11.3subsubsection.a.J.11.4subsubsection.a.J.11.5subsection.a.J.12subsubsection.a.J.12.1subsubsection.a.J.12.2subsubsection.a.J.12.3subsubsection.a.J.12.4subsubsection.a.J.12.5subsection.a.J.13subsubsection.a.J.13.1subsubsection.a.J.13.2subsubsection.a.J.13.3subsubsection.a.J.13.4subsubsection.a.J.13.5subsection.a.J.14subsubsection.a.J.14.1subsubsection.a.J.14.2subsubsection.a.J.14.3subsubsection.a.J.14.4subsection.a.J.15subsubsection.a.J.15.1subsubsection.a.J.15.2subsubsection.a.J.15.3subsubsection.a.J.15.4subsubsection.a.J.15.5subsection.a.J.16subsubsection.a.J.16.1subsubsection.a.J.16.2subsubsection.a.J.16.3subsubsection.a.J.16.4subsubsection.a.J.16.5subsection.a.J.17subsubsection.a.J.17.1subsubsection.a.J.17.2subsubsection.a.J.17.3subsubsection.a.J.17.4subsubsection.a.J.17.5subsection.a.J.18subsubsection.a.J.18.1subsubsection.a.J.18.2subsubsection.a.J.18.3subsubsection.a.J.18.4subsubsection.a.J.18.5subsection.a.J.19subsubsection.a.J.19.1subsubsection.a.J.19.2subsubsection.a.J.19.3subsubsection.a.J.19.4subsubsection.a.J.19.5
@@ Appendix A Summary of Pre-Registered Comparisons Between Twins and Humans
4. Context Effects (cognitive domain, replicates known human bias, preference measure, test of rationality, different versions of question, behavioral intention) Can digital twins demonstrate the attraction effect [25] and the compromise effect [54]? Humans replicated the attraction effect, but twins did not. Neither humans nor digital twins exhibited a compromise effect.
5. Defaults Effects (cognitive domain, replicates known human bias, preference measure, test of rationality, pro-social, different versions of question, behavioral intention, social desirability) Can digital twins demonstrate default effects [31, 48]? Human participants showed a default effect in a green energy adoption paradigm but not in an organ donation paradigm. Twins predicted a strong default effect in the organ donation paradigm but no effect in the green energy paradigm.
@@
7. Fees Accuracy (social domain, preference measure, different versions of question, attitude, scale question) (i) How does the knowledge of digital twins regarding fees and surcharges in the marketplace compare to their human counterparts? (ii) How do perceptions of fairness regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (iii) How does the variation in perceptions of fairness regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (iv) How do beliefs about government regulation regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (v) How does the variation in beliefs about government regulation regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (i) Twins were much more accurate than humans on knowledge questions. (ii) Twins closely mirrored the average human response on fairness questions. (iii) Twins exhibited narrower variance than humans on fairness perceptions. (iv) Twins were on average more favorable to government regulation compared to humans. (v) For both groups, conservatism was correlated with less support for pricing regulation. However, twins stereotyped participants based on their political ideology, with exaggerated support for regulation for liberal participants.
@@
9. Hiring algorithms (social domain, preference measure, human-tech interactions, different versions of question, scale question) Can digital twins predict job candidate preferences for workplace hiring policies? Specifically: (i) Can digital twins predict workplace preferences? (ii) Can experimenting on a population of digital twins recover the true average treatment effects as experimenting on a human population? (i) Twins tended to express views that were overall more pro-social (sustainability, ESG, work-life balance, transparent leadership, culture and values), and less "career-oriented" (firm prestige, career development) than those expressed by humans. (ii) Twins were overall less averse to hiring algorithms compared to humans.
@@
13. Obedient Twins (social domain, preference measure, valenced, scale question) LLMs are trained to be obedient and deferential. Does this tendency make digital twins more sensitive to survey instructions than their human counterparts? We tested this in three tasks: (i) Self-persuasion [8]. When prompted to consider the other side, do digital twins abandon their attitudes more readily than their human counterparts? (ii) Scenarios. Do digital twins more “obediently” follow the instructions to imagine themselves in different scenarios, leading to more sensitivity to scenario manipulations? (iii) Absurd Scenarios. Do digital twins “earnestly” respond to instructions that would be non-sensical to their human counterparts? (i) No. Twins were less sensitive to considering a different point of view compared to humans. (ii) No. Twins were less sensitive overall to these scenario manipulations compared to humans. (iii) No. Humans were more sensitive to the instructions to forecast feelings in absurd scenarios.
@@
15. Preferences for Redistribution (social domain, attitude, scale question) Can digital twins predict preferences for redistribution and other outcomes like trust, socio-economic background, and beliefs on fairness and social mobility? Twins showed more extreme redistribution preferences. On average, they were more likely to think "people should take care of themselves" but more willing to pay taxes to improve healthcare for all people. They were more likely to think people are fair and that people can be trusted. They were also more likely to think that hard work and luck are equally important for success. Finally, they showed very little distribution in their fathers’ educational attainment, with almost 100% of them reporting their father only completed high school.
@@ Appendix C Details for Persona Construction
In this section, we detail the construction of the full persona, persona summary, and demographics persona versions of digital twins.
@@ Appendix D Distribution of Each Performance Metric Across Outcomes
Figure 9: Distribution Across Outcomes of |Glass’s $\Delta$| Between Human and Twin Responses.
@@
Appendix E Meta-analysis of Other Performance Metrics
@@
Figure 12: Results from Meta-Analysis (Mixed Linear Model with |Glass’s $\Delta$| as Dependent Variable).
@@ Appendix G Comparison Across Implementation Methods
G.1 Alternative base LLMs
We test GPT-5, Deepseek (deepseek-r1-0528), Gemini (gemini-2.5-flash), and a fine-tuned GPT-4.1 as alternative base models. Full persona prompts were used by all benchmarks. The temperature for GPT-4.1 and its fine-tuned variants was set to the default value of 0.7. For DeepSeek (deepseek-r1-0528) and Gemini (gemini-2.5-flash), we used a temperature of 0, as we observed that this setting improves correlation with the full persona. The temperature for GPT-5 was fixed at 1 due to the API restriction. All other API parameters for these benchmarks were kept at their default values.
@@ G.1 Alternative base LLMs
G.2 Alternative Temperature
We replicate the main comparison in the paper with a temperature of 0 instead of the default temperature. Figure 16 displays the results, including the full persona with default temperature as baseline. We see that when temperature is 0, like when temperature is set to default, digital twins based on full personas perform significantly better than generic twins based on demographics on correlation, and not significantly differently on individual-level accuracy. We also see that performance is overall improved when temperature=0.
@@ Appendix H Centaur Experiments
The state-of-the-art LLMs we used as based models are general-purpose models (e.g., GPT-4.1). A natural question arises: what if an LLM fine-tuned specifically for human behavioral prediction was used as base model for the creation of digital twins? One such potential base model is Centaur [6], a 70B-parameter model fine-tuned from Llama-3.1-70B using data from over 60,000 participants, encompassing 10 million choices across 160 behavioral experiments.
@@
To formally test the difference between human and digital twin prediction accuracy, we estimated a linear mixed-effects model: err_it = β_0 + β_1Human_it + β_2Task_it + (1 | ResponseId_i) + ϵ_it where $err_{it}$ is the prediction error for respondent $i$ on task $t$ (abs(predicted value in treatment condition - observed value in treatment condition)/(observed value in treatment condition)), $Human_{it}$ indicates predictions for human vs. digital twins, $Task_{it}$ indexes the five tasks, $(1|ResponseId_{i})$ is a random intercept for each respondent, and $\epsilon_{it}$ is the residual. Results confirmed significantly lower errors for predictions of human responses on average ($\beta$ = -0.072, SE = 0.011, p << 0.001) (see Table 2).
@@
The spread of misinformation across social media is a growing public policy concern, with severe implications for health, politics, and science. Recent work suggests that people care about being accurate but share misinformation because the social media context focuses their attention on other factors [46]. The proposed solution for inattention-based misinformation sharing is to prime accuracy before the decision to share information. Research testing this account [43] demonstrates that asking people to rate the accuracy of unrelated headlines or state the extent to which they agree that it is important to share accurate content on social media before sharing improves truth discernment: that is, reduces sharing of false headlines compared to true headlines. The current study aims to replicate this pattern with human participants, and test whether their digital twins respond similarly to these accuracy nudges.
Specifically, this study aims to replicate study 5 of [43], using a different set of headlines. While [43] demonstrate the effectiveness of accuracy nudges in the context of political news headlines, the current study utilizes their paradigm to test the effectiveness of accuracy nudges in the context of non-political entertainment news headlines. The decision to sharing entertaining news may provoke consideration of how entertaining the headline is, beyond considerations on the accuracy of the news and trustworthiness of the sources [34]. Indeed, people generally prefer and are most likely to engage with entertaining news [23, 49, 62] and interesting or emotionally arousing word-of-mouth [4, 3, 5]. Testing the effectiveness of accuracy nudges for entertaining headlines is therefore an important extension of the prior work.
@@
A total of 1,003 human participants on Prolific (and their digital twins) took part in the study (50% female, 50% male, 0% other; $M_{\text{age}}=49.37$, $SD=14.55$), following the protocol proposed by [43]. Participants were randomly assigned to the same condition in a $2$ (Headline Veracity: False vs. True; within-subjects) $\times$ 44 (Intervention: Control vs. Active Control vs. Treatment vs. Importance Treatment; between-subjects) mixed design. All participants were asked to indicate how likely they would be to share 16 headlines, displayed in random order, as the dependent variable. Participants in the Control condition proceeded directly to this task, while participants in the Active Control, Treatment, and Importance Treatment conditions were first asked to complete the following separate tasks:
Active Control: Participants were first told that they would help pretest actual news headlines for future studies. They were then shown a headline and asked to rate, “In your opinion, is the above headline funny, amusing, or entertaining?” (1-Extremely unfunny, 6-Extremely funny). They were randomly assigned to see one of four headlines—two of which were true and two of which were false. The stimuli were the same as [43], with the exception that images were not included, given that the digital twins cannot process visual images.
Treatment: Similarly, participants were told that they would help pretest actual news headlines for future studies. They were then shown two headlines in randomized order and asked to rate, “In your opinion, is the above headline accurate?” (1-Extremely inaccurate, 6-Extremely accurate). They were randomly assigned to see one of four headlines—two of which were true and two of which were false. The stimuli were the same as [43], again with the exception that images were not included.
@@
All participants then responded to the dependent variable. Participants were presented with 16 news headlines in randomized order. For each headline, participants were asked to indicate, “If you were to see the above article on social media, how likely would you be to share it?” (1-Extremely unlikely, 6-Extremely likely). All headlines were pretested to be entertaining and interesting, but not identity relevant [34]. Eight of the headlines were inspired by real headlines from the websites of trustworthy news organizations. These headlines were paired with generally trustworthy news sources. The remaining 8 headlines were false headlines. These headlines were paired with generally untrustworthy news sources. Source trustworthiness perceptions were confirmed at the end of the survey. Table LABEL:SI_accuracy_tab1 provides the headline and source pairings.
@@
Participants in the Control, Active Control, and Treatment condition next responded to the same question shown earlier to participants in the Importance Treatment condition: “Do you agree or disagree that ‘it is important to only share content on social media that is accurate and unbiased’?” (1-Strongly agree, 6-Strongly disagree). All participants then responded to exploratory questions intended to capture stated social media sharing preferences. As a measure of goal prioritization when sharing entertaining news, participants were first asked to indicate, “When you share information with others on social media, is it more important that the information is:” (1-Verifiably correct, 6-Entertaining; [34]). Participants then indicated, “Would you ever consider sharing news like you saw in the study today on social media?”.
@@
A linear mixed-effects regression was estimated [^15] Linear mixed-effects models were estimated using the lmerTest package, which provides $p$-values and $F$-tests using Satterthwaite’s approximation. with participant as a random effect, average sharing intentions as the dependent variable, and Participant Type (Human =0=0, Digital Twin =1=1), Headline Veracity (False =0=0, True =1=1), Intervention (Control =0=0, Active Control =1=1, Treatment =2=2, Importance Treatment =3=3), and their interactions as predictors. The results revealed a significant main effect of Participant ($F(1,2997)=33.41$, $p<.001$), a non-significant main effect of Headline Veracity ($F(1,2997)=1.24$, $p=.27$), along with a significant main effect of Intervention ($F(3,2997)=8.89$, $p<.001$), qualified by a significant three-way interaction ($F(3,2997)=5.75$, $p<.001$), indicating that, at an aggregate level, accuracy nudges affected how the digital twins shared true vs. false headlines differently than humans. Figure 19 shows average sharing intentions across all conditions for humans (Panel A) and their digital twins (Panel B). See Table LABEL:SI_accuracy_tab2 for the full regression results.

### v4

@@ Discussion
We hope that by making our data and code publicly available and by providing a detailed analysis of each study (see SI), we will help the field develop a data-driven, impartial understanding of the current opportunities and challenges presented by digital twins.
Acknowledgments
Caroline Weaver provided research assistance on the Context Effects and Defaults Effects substudies as well as the design, implementation and analysis of the survey reported in SI.I. Xuwen Hua provided research assistance on the Context Effects and Defaults Effects substudies. Arnab Choudhury provided research assistance on the Hiring Algorithms substudy. Sanjana Rosario provided research assistance on the Quantitative Intuition substudy.
Author contributions:
G.G., D.J.M., M.B, E.J.J., V.M., A.A., S.B., D.D., H.F., E.F., A.G., M.H., K.J., B.K., A.K., K.L., H.L., P.P., O.N., O.T. designed the substudies, analyzed the data from each substudy and wrote the corresponding sections in SI.J; O.T. ran the studies on human participants; T.P., G.G., G.J.F., M.B.S., O.T. ran the studies on digital twins; T.P., G.G., D.J.M., G.J.F., M.B.S., O.T. produced the analyses reported in the main paper and SI.D-H; E.J.J., O.T. designed and analyzed the study reported in SI.I; M.B., E.J.J., O.T. were particularly involved in writing the main paper, with input from multiple co-authors.
Competing interests:
The authors declare no competing interests.
Data and materials availability:
Data are publicly available at https://huggingface.co/datasets/LLM-Digital-Twin/Twin-2K-500-Mega-Study. The only piece of data which is not publicly available is the Prolific ID of the participants. All coded is publicly available at https://github.com/TianyiPeng/Twin-2K-500-Mega-Study.
Materials and Methods
Each sub-study was first run on Prolific according to its pre-registered plan (https://researchbox.org/4145). All sub-studies were run from the same Prolific account to ensure consistency. Invitations were sent to Prolific participants from the Twin-2K-500 panel. Each sub-study was approved by Columbia University’s IRB. All sub-studies were run from April to June 2025.
Next, each sub-study was conducted on the digital twins of its participants. To set up the simulation, we extracted twin information from the Twin-2K-500 dataset and provided it as prompt input. For each human participant, the completed survey was transformed into text, stripped of its answers, and incorporated into the user prompt. Survey questions varied across participants due to randomization; we preserved these differences when generating inputs for their twins. The LLM was then tasked with producing responses to these surveys. The creation of digital twins and the related computations were approved by Columbia University’s IRB under Protocol IRB-AAAV5832.
For each persona in each sub-study, we constructed a user prompt using a template (see SI) and queried the LLM via its API (OpenAI API for OpenAI models, OpenRouter API for others). The JSON outputs were then post-processed to ensure all questions were answered in the required format (e.g., no missing values or out-of-range responses). If validation failed, the API call was retried. A complete version of the prompt template, including placeholders for persona profiles, new questions, and JSON-based response formatting, is provided in the SI.
@@ Acknowledgments
Digital Twins as Funhouse Mirrors: Five Key Distortions
@@ Appendix A Summary of Pre-Registered Comparisons Between Twins and Humans
4. Context Effects (cognitive domain, replicates known human bias, preference measure, test of rationality, different versions of question, behavioral intention) Can digital twins demonstrate the attraction effect (?) and the compromise effect (?)? Humans replicated the attraction effect, but twins did not. Neither humans nor digital twins exhibited a compromise effect.
5. Defaults Effects (cognitive domain, replicates known human bias, preference measure, test of rationality, pro-social, different versions of question, behavioral intention, social desirability) Can digital twins demonstrate default effects (?, ?)? Human participants showed a default effect in a green energy adoption paradigm but not in an organ donation paradigm. Twins predicted a strong default effect in the organ donation paradigm but no effect in the green energy paradigm.
@@
7. Fees Accuracy (social domain, preference measure, different versions of question, attitude, scale question) (i) How does the knowledge of digital twins regarding fees and surcharges in the marketplace compare to that of their human counterparts? (ii) How do perceptions of fairness regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (iii) How does the variation in perceptions of fairness regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (iv) How do beliefs about government regulation regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (v) How does the variation in beliefs about government regulation regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts? (i) Twins were much more accurate than humans on knowledge questions. (ii) Twins closely mirrored the average human response on fairness questions. (iii) Twins exhibited narrower variance than humans on fairness perceptions. (iv) Twins were on average more favorable to government regulation compared to humans. (v) For both groups, conservatism was correlated with less support for pricing regulation. However, twins stereotyped participants based on their political ideology, with exaggerated support for regulation for liberal participants.
@@
9. Hiring algorithms (social domain, preference measure, human-tech interactions, different versions of question, scale question) Can digital twins predict job candidate preferences for workplace hiring policies? Specifically: (i) Can digital twins predict workplace preferences? (ii) Can experimenting on a population of digital twins recover the true average treatment effects as experimenting on a human population? (i) Twins tended to express views that were overall more pro-social (sustainability, ESG, work-life balance, transparent leadership, culture and values), and less ”career-oriented” (firm prestige, career development) than those expressed by humans. (ii) Twins were overall less averse to hiring algorithms compared to humans.
@@
13. Obedient Twins (social domain, preference measure, valenced, scale question) LLMs are trained to be obedient and deferential. Does this tendency make digital twins more sensitive to survey instructions than their human counterparts? We tested this in three tasks: (i) Self-persuasion (?). When prompted to consider the other side, do digital twins abandon their attitudes more readily than their human counterparts? (ii) Scenarios. Do digital twins more “obediently” follow the instructions to imagine themselves in different scenarios, leading to more sensitivity to scenario manipulations? (iii) Absurd Scenarios. Do digital twins “earnestly” respond to instructions that would be non-sensical to their human counterparts? (i) No. Twins were less sensitive to considering a different point of view compared to humans. (ii) No. Twins were less sensitive overall to these scenario manipulations compared to humans. (iii) No. Humans were more sensitive to the instructions to forecast feelings in absurd scenarios.
@@
15. Preferences for Redistribution (social domain, attitude, scale question) Can digital twins predict preferences for redistribution and other outcomes like trust, socio-economic background, and beliefs on fairness and social mobility? Twins showed more extreme redistribution preferences. On average, they were more likely to think ”people should take care of themselves” but more willing to pay taxes to improve healthcare for all people. They were more likely to think people are fair and that people can be trusted. They were also more likely to think that hard work and luck are equally important for success. Finally, they showed very little distribution in their fathers’ educational attainment, with almost 100% of them reporting their father only completed high school.
@@ Appendix C Details for Persona Construction
In this section, we detail the construction of the various types of personas. That is, we provide the text that is inserted into the full prompt template (Section B.1), as {Persona Profile}.
@@ C.3 Demographics Persona
C.4 Empty Persona
The Empty Persona specification does not provide any individual-level information in the prompts. Instead, {Persona Profile} is replaced with the placeholder [Empty Persona Profile].
Specifically, the prompt used for all individuals is:
Persona Profile (This individual’s past survey responses):
[Empty Persona Profile]
@@ Appendix D Distribution of Each Performance Metric Across Outcomes
Figure 9: Distribution Across Outcomes of —Glass’s $\Delta$— Between Human and Twin Responses.
@@
Appendix E Digital Twin Performance Across Domains
We explore patterns in performance across domains using a meta-analytic approach to understand where twins are more likely to fall short. Given that correlation is the performance metric most sensitive to the individual-level information captured by digital twins, we focus our discussion on the meta-analysis regression with the (z-transformed) correlation for each outcome as the dependent variable, and report the meta-analysis of the other performance metrics for completeness. [^11] We z-transform the correlation due to better statistical properties, e.g., being approximately normally distributed. Our independent variables include a set of categorical labels that characterize each outcome, related to the following domains: social, preferences/attitudes, cognitive skills/rationality, content evaluation, human-tech interactions. We also include mechanical features that might influence performance (e.g., sample size).
The categorical labels in our meta-analysis are not mutually exclusive; an outcome may be tagged with multiple labels. All labels are binary, except sample size. The labels are as follows (with examples from our data):
  • Social
    • Social domain (104 outcomes): involves social topics (e.g., likelihood of sharing news content with others)
    • Social cognition (15 outcomes): involves social cognition such as asking the participant to judge the behavior of others (e.g., judging someone who donates to both political parties)
  • Preferences/attitudes
  • Cognitive skills/rationality
    • Test of rationality (10 outcomes): assesses rational decision-making (e.g., sensitivity to context effects)
  • Content evaluation
    • Valenced (40 outcomes): involves positive versus negative judgments (e.g., judging the creativity of ideas)
  • Human–tech interactions
  • Mechanical
  • Digital twins performed relatively better in the social domain. The correlation was higher when outcomes related to conflict, pro-social issues, social cognition, or personality. However, the correlation was significantly lower for outcomes where social desirability was salient, suggesting twins are less capable of mimicking human responses in socially-sensitive contexts (descriptive evidence suggests twins are more likely to provide socially-desirable responses).
  • The labels related to preferences and attitudes were not significant, i.e., correlation did not differ from average, except for outcomes related to political preferences: correlations were particularly low for outcomes in the political domain. This is consistent with previous findings such as (?) or (?).
  • Correlation was on average higher in the cognitive domain, although further analysis reveals that digital twins tend to systematically deviate from humans in this domain, by being more rational and knowledgeable.
  • Twins did not correlate with humans better or worse on average when evaluating content (i.e., the coefficient associated with content evaluation was not significant in the meta-analysis regressions). However, their responses were less correlated with humans when the evaluation was valenced (positive versus negative).
  • When questions varied across participants, twin-human correlation dropped, consistent with prior findings that LLMs struggle to treat prompt-level variation as exogenous (?).
  • Twins showed higher correlation with humans on outcomes that use response scales. This may be due to the fact that a significant proportion of their training data comes from questions with a similar format.
Figure 11: Results from Meta-Analysis (Mixed Linear Model with z-transformed Correlation as Dependent Variable). Regression coefficients with 95% confidence intervals. The average correlation across all outcomes is 0.197. Correlation between responses from twins versus their human counterparts tends to be higher in social domains, except when social desirability is salient. Correlation tends to be higher in the cognitive domain and in domains related to human-technology interactions, but lower in the political domain and when providing valenced evaluations.
@@ Appendix E Digital Twin Performance Across Domains
Figure 13: Results from Meta-Analysis (Mixed Linear Model with —Glass’s $\Delta$— as Dependent Variable).
@@ Appendix G Comparison Across Implementation Methods
G.1 Persona summaries
We add as a benchmark an alternative way to construct digital twins, labeled “persona summary:” a concise, statement-based summary (approximately 13K characters) of the questions and responses with distributional information (e.g., Big 5 personality scores with percentile ranks rather than 44 detailed questions and answers on which the scores are based). Results are reported in Figure 16. We see that this version performs very similarly to the full persona, offering a viable lower-cost alternative.
Figure 16: Gains from Leveraging Individual-Level Data, Including Persona Summary as Benchmark. *: best performing benchmark, or not significantly different from best at p<<0.05 (not applicable to ratio of standard deviations).
G.2 Alternative base LLMs
We test GPT-5, Deepseek (deepseek-r1-0528), Gemini (gemini-2.5-flash, gemini-3-pro), and a fine-tuned GPT-4.1 as alternative base models. Full persona prompts were used by all benchmarks. The temperature for GPT-4.1 and its fine-tuned variants was set to the default value of 0.7. For the other models we selected the temperature that was expected to yield the best performance. We used a temperature of 0 for DeepSeek (deepseek-r1) and Gemini-2.5 (gemini-2.5-flash) because we found this helped GPT4.1. We used the default temperature of 1 for Gemini-3-pro, as Google recommends not changing the temperature of this model for better performance. We set the temperature for GPT-5 at 1 due to the API restriction. All other API parameters for these benchmarks were kept at their default values.
@@ G.2 Alternative base LLMs
G.3 Alternative Temperature
We replicate the main comparison in the paper (i.e., with GPT-4.1) with a temperature of 0 instead of the default temperature. Figure 18 displays the results, including the full persona with default temperature as baseline. We see that when temperature is 0, like when temperature is set to default, digital twins based on full personas perform significantly better than generic twins based on demographics on correlation, and not significantly differently on individual-level accuracy. We also see that performance is overall improved when temperature=0.
@@ Appendix H Centaur Experiments
The state-of-the-art LLMs we used as based models are general-purpose models (e.g., GPT-4.1). A natural question arises: what if an LLM fine-tuned specifically for human behavioral prediction was used as base model for the creation of digital twins? One such potential base model is Centaur (?), a 70B-parameter model fine-tuned from Llama-3.1-70B using data from over 60,000 participants, encompassing 10 million choices across 160 behavioral experiments.
@@
To formally test the difference between human and digital twin prediction accuracy, we estimated a linear mixed-effects model:
$err_{it}=\beta_{0}+\beta_{1}Human_{it}+\beta_{2}Task_{it}+(1|ResponseId_{i})+\epsilon_{it}$
where $err_{it}$ is the prediction error for respondent $i$ on task $t$ (abs(predicted value in treatment condition - observed value in treatment condition)/(observed value in treatment condition)), $Human_{it}$ indicates predictions for human vs. digital twins, $Task_{it}$ indexes the five tasks, $(1|ResponseId_{i})$ is a random intercept for each respondent, and $\epsilon_{it}$ is the residual. Results confirmed significantly lower errors for predictions of human responses on average ($\beta$ = -0.072, SE = 0.011, p << 0.001) (see Table 2).
@@
The spread of misinformation across social media is a growing public policy concern, with severe implications for health, politics, and science. Recent work suggests that people care about being accurate but share misinformation because the social media context focuses their attention on other factors (?). The proposed solution for inattention-based misinformation sharing is to prime accuracy before the decision to share information. Research testing this account (?) demonstrates that asking people to rate the accuracy of unrelated headlines or state the extent to which they agree that it is important to share accurate content on social media before sharing improves truth discernment: that is, reduces sharing of false headlines compared to true headlines. The current study aims to replicate this pattern with human participants, and test whether their digital twins respond similarly to these accuracy nudges.
Specifically, this study aims to replicate study 5 of (?), using a different set of headlines. While (?) demonstrate the effectiveness of accuracy nudges in the context of political news headlines, the current study utilizes their paradigm to test the effectiveness of accuracy nudges in the context of non-political entertainment news headlines. The decision to sharing entertaining news may provoke consideration of how entertaining the headline is, beyond considerations on the accuracy of the news and trustworthiness of the sources (?). Indeed, people generally prefer and are most likely to engage with entertaining news (?, ?, ?) and interesting or emotionally arousing word-of-mouth (?, ?, ?). Testing the effectiveness of accuracy nudges for entertaining headlines is therefore an important extension of the prior work.
@@
A total of 1,003 human participants on Prolific (and their digital twins) took part in the study (50% female, 50% male, 0% other; $M_{\text{age}}=49.37$, $SD=14.55$), following the protocol proposed by (?). Participants were randomly assigned to one condition in a $2$ (Headline Veracity: False vs. True; within-subjects) $\times$ 44 (Intervention: Control vs. Active Control vs. Treatment vs. Importance Treatment; between-subjects) mixed design (each human and their digital twin were assigned to the same condition). All participants were asked to indicate how likely they would be to share 16 headlines, displayed in random order, as the dependent variable. Participants in the Control condition proceeded directly to this task, while participants in the Active Control, Treatment, and Importance Treatment conditions were first asked to complete the following separate tasks:
Active Control: Participants were first told that they would help pretest actual news headlines for future studies. They were then shown a headline and asked to rate, “In your opinion, is the above headline funny, amusing, or entertaining?” (1-Extremely unfunny, 6-Extremely funny). They were randomly assigned to see one of four headlines—two of which were true and two of which were false. The stimuli were the same as (?), with the exception that images were not included, given that the digital twins cannot process visual images.
Treatment: Similarly, participants were told that they would help pretest actual news headlines for future studies. They were then shown two headlines in randomized order and asked to rate, “In your opinion, is the above headline accurate?” (1-Extremely inaccurate, 6-Extremely accurate). They were randomly assigned to see one of four headlines—two of which were true and two of which were false. The stimuli were the same as (?), again with the exception that images were not included.
@@
All participants then responded to the dependent variable. Participants were presented with 16 news headlines in randomized order. For each headline, participants were asked to indicate, “If you were to see the above article on social media, how likely would you be to share it?” (1-Extremely unlikely, 6-Extremely likely). All headlines were pretested to be entertaining and interesting, but not identity relevant (?). Eight of the headlines were inspired by real headlines from the websites of trustworthy news organizations. These headlines were paired with generally trustworthy news sources. The remaining 8 headlines were false headlines. These headlines were paired with generally untrustworthy news sources. Source trustworthiness perceptions were confirmed at the end of the survey. Table LABEL:SI_accuracy_tab1 provides the headline and source pairings.
@@
Participants in the Control, Active Control, and Treatment condition next responded to the same question shown earlier to participants in the Importance Treatment condition: “Do you agree or disagree that ‘it is important to only share content on social media that is accurate and unbiased’?” (1-Strongly agree, 6-Strongly disagree). All participants then responded to exploratory questions intended to capture stated social media sharing preferences. As a measure of goal prioritization when sharing entertaining news, participants were first asked to indicate, “When you share information with others on social media, is it more important that the information is:” (1-Verifiably correct, 6-Entertaining; (?)). Participants then indicated, “Would you ever consider sharing news like you saw in the study today on social media?”.
@@
A linear mixed-effects regression was estimated [^13] Linear mixed-effects models were estimated using the lmerTest package, which provides $p$-values and $F$-tests using Satterthwaite’s approximation. with participant as a random effect, average sharing intentions as the dependent variable, and Participant Type (Human =0=0, Digital Twin =1=1), Headline Veracity (False =0=0, True =1=1), Intervention (Control =0=0, Active Control =1=1, Treatment =2=2, Importance Treatment =3=3), and their interactions as predictors. The results revealed a significant main effect of Participant Type ($F(1,2997)=33.41$, $p<.001$), a non-significant main effect of Headline Veracity ($F(1,2997)=1.24$, $p=.27$), along with a significant main effect of Intervention ($F(3,2997)=8.89$, $p<.001$), qualified by a significant three-way interaction ($F(3,2997)=5.75$, $p<.001$), indicating that, at an aggregate level, accuracy nudges affected how the digital twins shared true vs. false headlines differently than humans. Figure 21 shows average sharing intentions across all conditions for humans (Panel A) and their digital twins (Panel B). See Table LABEL:SI_accuracy_tab2 for the full regression results.